OpenStudy (anonymous):

Graph 4x^2 + y^2 = 9. What are its lines of symmetry?

OpenStudy (mertsj):

Do you recognize that this is an ellipse?

OpenStudy (anonymous):

no I did not please explain.. :)

OpenStudy (mertsj):

This is one of those situations that arise in mathematics when you have to recognize when there is something to reason out and when there is something to learn. This time there is something to learn. You must learn to recognize the conic sections by looking at their equations.

OpenStudy (mertsj):

I will try to help you.

OpenStudy (anonymous):

Ok would you like the multiple choice answers?

OpenStudy (mertsj):

In a circle, you must have x^2 and y^2. The signs must be positive and the coefficients must be the same.

OpenStudy (mertsj):

In a parabola, you must have x and y^2 or y and x^2. Then you know the graph is a parabola.

OpenStudy (mertsj):

In an ellipse you must have x^2 and y^2, both positive but the coefficients are not the same.

OpenStudy (mertsj):

In a hyperbola, you must have x^2 and y^2 but one of them is negative.

OpenStudy (mertsj):

So now we know that 4x^2 + y^2 = 9 is an ellipse because we have x^2 and y^2. Both are positive and the coefficients are different.

OpenStudy (anonymous):

ok.

OpenStudy (mertsj):

So now your job is to rearrange it so it is in this form: \[\frac{x^2}{a^2}+\frac{y^2}{b^2}=1\]

OpenStudy (mertsj):

Let me know when you have that done.

OpenStudy (anonymous):

umm ok well im not completely sure of how to rearange i t

OpenStudy (mertsj):

Do you notice that the right side has to be 1?

OpenStudy (anonymous):

ah yes

OpenStudy (mertsj):

What is the right side of your equation?

OpenStudy (anonymous):

the one?

OpenStudy (mertsj):

Here is your problem: 4x^2 + y^2 = 9 What is the right side?

OpenStudy (anonymous):

ok the 9

OpenStudy (mertsj):

But it is supposed to be 1. So you must answer the question, "What should I do to the 9 to turn it into a 1?"

OpenStudy (anonymous):

Divide by 9? Im just a little confused.

OpenStudy (mertsj):

Yes. Excellent!! Divide by 9. Divide EVERY term on BOTH sides by 9. Post your result.

OpenStudy (anonymous):

4x^2 + y^2 = 9 ok so 4x^2/9 + y^2/9 = 1?

OpenStudy (mertsj):

Yes. Perfect. Now remember, we want it to be in the form: \[\frac{x^2}{a^2}+\frac{y^2}{b^2}=1\] and we have it in the form: \[\frac{4x^2}{9}+\frac{y^2}{9}=1\]

OpenStudy (mertsj):

So we are not quite there.

OpenStudy (mertsj):

Do you know that multiplying by 4/9 is the same as dividing by 9/4?

OpenStudy (anonymous):

so it would be 4x^2 y^2 ------ + ----------- 9x^2 9^2

OpenStudy (mertsj):